# QoS-Aware Load Balancing in the Computing Continuum via Multi-Player Bandits

### v1

@@ Abstract
As computation shifts from the cloud to the edge to reduce processing latency and network traffic, the resulting Computing Continuum (CC) creates a dynamic environment where it is challenging to meet strict Quality of Service (QoS) requirements and avoid service instance overload. Existing methods often prioritize global metrics, overlooking per-client QoS, which is crucial for latency-sensitive and reliability-critical applications. We propose QEdgeProxy, a decentralized QoS-aware load balancer that acts as a proxy between IoT devices and service instances in CC. We formulate the load balancing problem as a Multi-Player Multi-Armed Bandit (MP-MAB) with heterogeneous rewards, where each load balancer autonomously selects service instances that maximize the probability of meeting its clients’ QoS targets by using Kernel Density Estimation (KDE) to estimate QoS success probabilities. It also incorporates an adaptive exploration mechanism to recover rapidly from performance shifts and non-stationary conditions. We present a Kubernetes-native QEdgeProxy implementation and evaluate it on an emulated CC testbed deployed on a K3s cluster with realistic network conditions and a latency-sensitive edge-AI workload. Results show that QEdgeProxy significantly outperforms proximity-based and reinforcement-learning baselines in per-client QoS satisfaction, while adapting effectively to load surges and instance availability changes.
@@ I Introduction
The growing scale of Internet of Things (IoT) deployments has made routing all data to remote cloud servers for processing increasingly inefficient. This centralized model leads to increased latency, congestion, and limited responsiveness for real-time applications. The Computing Continuum (CC) addresses these challenges by distributing computation across IoT, edge, fog, and cloud nodes, enabling data to be processed closer to its source to improve both performance and resource utilization.
Despite these advantages, the CC remains highly dynamic: nodes vary in load, availability, and network conditions, while IoT workloads fluctuate in time and space. Ensuring that services deployed across this heterogeneous continuum continue to meet strict Quality of Service (QoS) requirements therefore demands mechanisms that can adapt to changing conditions.
@@
Reliable service delivery in the CC depends on the explicit definition and enforcement of QoS requirements. These requirements specify measurable performance guarantees, typically expressed through Service Level Objectives (SLOs) and Service Level Agreements (SLAs) [1]. SLOs define concrete targets—–such as ensuring that a given percentage of requests remain below a latency threshold—–while SLAs formalize these guarantees contractually between providers and consumers. QoS requirements in the CC may be expressed along multiple dimensions, including latency, throughput, reliability, availability, privacy, and security. Their importance varies across application domains; for example, industrial control and autonomous systems prioritize ultra-low latency, whereas healthcare applications may emphasize both privacy and availability. Formalizing such QoS constraints provides a foundation for coordinating placement, routing, and adaptation strategies so that services remain within their expected performance boundaries under dynamic and heterogeneous CC conditions.
@@
This work focuses on two key challenges that arise once service instances are deployed in the CC: (i) guaranteeing QoS through adaptive, decentralized request routing under uncertainty, and (ii) continuously delivering data across different layers of the continuum without overloading service instances on resource-constrained devices or compute nodes. To support QoS-aware routing in dynamic CC conditions, in our previous work [2], we introduced QEdgeProxy, a lightweight decentralized load balancer deployed on every CC node. QEdgeProxy acts as an intermediary between clients and service instances, selecting suitable instances based on locally observed QoS and continuously adapting its routing choices as conditions change. Rather than selecting the instance that maximizes the QoS, it maintains a pool of instances capable of meeting the QoS target with high probability, named QoS pool, enabling effective load balancing while preserving per-client QoS guarantees. QEdgeProxy is designed to handle both client-to-service and service-to-service communication, allowing IoT devices and intermediate services to forward requests through the proxy deployed on their nearest or hosting node. By maintaining and dynamically updating a QoS-feasible pool of service instances, QEdgeProxy ensures stable request distribution and mitigates overload while satisfying QoS constraints across the CC.
@@
Building on this prior work, where QEdgeProxy was introduced together with a high-level architectural design and a proof-of-concept implementation, the load-balancing logic in [2] remained limited to constructing QoS pools from the average of latency measurements and distributing requests within the pool using simple round-robin scheduling. In this paper, we significantly extend and formalize the load-balancing mechanism by developing the underlying algorithms for QoS-pool maintenance and request routing. We formulate the problem as a Multi-Player Multi-Armed Bandit (MP-MAB) with heterogeneous rewards and provide a decentralized learning procedure—built on KDE-based QoS estimation, adaptive exploration, and instance-specific weight updates—that enables QEdgeProxy to operate robustly under dynamic and non-stationary CC conditions. This work therefore provides the theoretical foundation, the detailed algorithmic formulation, and a comprehensive evaluation that were not addressed in the original QEdgeProxy prototype, resulting in a substantially more complete and rigorous load-balancing solution.
@@
  • Decentralized QoS-aware load balancing via MP-MAB. We introduce a fully decentralized set of algorithms for QoS-aware load balancing, formulated as a MP-MAB problem with heterogeneous and non-stationary rewards. Each node autonomously maintains a QoS pool—a set of service instances predicted to satisfy the required QoS—using a combination of $\varepsilon$-decay exploration and Kernel Density Estimation (KDE) to estimate instance-level QoS success probabilities. Furthermore, it incorporates an adaptive exploration mechanism that reduces exploration as confidence grows but increases it when QoS degradation or performance shifts occur, enabling rapid recovery in non-stationary conditions. Requests are then routed using a Smooth Weighted Round Robin (SWRR) mechanism parameterized by these probabilities, enabling balanced utilization of service instances while preserving per-client latency guarantees.
@@
  • Kubernetes-native implementation. We implement the algorithms as a Kubernetes-native component, enabling each node to run a decentralized QoS-aware load balancer that maintains QoS pools, performs adaptive weighted routing, and obtains node and instance updates through an MQTT-based informer to reduce the monitoring overhead. This design demonstrates that decentralized QoS-aware load balancing can be integrated into existing CC orchestration solutions with minimal configuration and without modifying the application code.
@@
  • Real-world validation in a distributed CC testbed. We evaluate our solution deployed across a real K3s cluster emulating a 30-node CC topology with realistic network latencies. Using a latency-sensitive IoT edge-AI workload, we show that our solution outperforms proximity-based and reinforcement-learning baselines, achieving higher per-client QoS satisfaction, improved fairness, robust adaptiveness to dynamic changes, and low resource overhead suitable for constrained edge devices.
@@
The remainder of this paper is organized as follows. Section II reviews related work on QoS-aware routing, decentralized load balancing, and learning-based decision-making in the CC. Section III introduces the two-phased system model, describing both design-time and runtime behavior of the system, and QoS requirements definition. Section IV formalizes decentralized QoS-aware load balancing as an MP-MAB problem with heterogeneous rewards. Section V proposes decentralized load-balancing algorithms as implementation of the QEdgeProxy, including QoS-pool construction, KDE-based estimation, $\varepsilon$-decay exploration, and SWRR-based request routing. Section VI presents the Kubernetes-native implementation of the QEdgeProxy, and Section VII reports the experimental evaluation conducted on an emulated CC testbed. Finally, Section VIII concludes the paper and discusses directions for future work. In Table I we listed all the symbols used in the paper.
@@
$\mathcal{K}$ Set of load balancers (one per CC node)
@@
$\mathcal{M}(t)$ Active service instances at time $t$
@@
$r_{k,m}(t)$ Reward for LB $k$ routing a request to $m$ at time $t$
@@
$\mu_{k,m}(t)$ Expected QoS success probability for LB $k$ on instance $m$ at time $t$
@@
$\varepsilon(t)$ Exploration factor at time $t$
@@ II Related work
Fahs and Pierre introduced Proxy-mity [3], a proximity-aware traffic routing system designed as a plugin for Kubernetes. Their goal was to make Kubernetes’ routing aware of geographical distribution by enabling a trade-off between load balancing and latency reduction through a control parameter $\alpha$. When $\alpha=0$, traffic is evenly distributed across all service replicas, whereas $\alpha=1$ prioritizes routing to the closest instance. The approach seamlessly integrates with Kubernetes and reduces the request latency by up to 90% in experimental fog setups. However, Proxy-mity considers only network latency as a proximity metric and does not incorporate feedback from actual request processing times. Moreover, it does not allow the explicit definition of QoS constraints such as maximum latency thresholds or success ratios.
Liu and Fang [4] proposed AdeQoS, a decentralized multi-player multi-armed bandit (MP-MAB) algorithm that guarantees per-player QoS satisfaction. Their model assumes that each player (e.g., a load balancer) interacts with a set of arms (service instances) without coordination or observation of others’ actions. Unlike conventional reward maximization, AdeQoS tries to ensure that each player maintains an expected reward above a given QoS threshold, achieving $O(1)$ QoS regret. This approach is closely aligned with our formulation (see Section IV), as both aim to meet per-client QoS constraints rather than maximizing aggregate utility. However, AdeQoS assumes that the number of arms exceeds the number of players and that collisions occur when two players select the same arm, which differs from our model (Section III) where each service instance can handle multiple requests simultaneously. Moreover, AdeQoS relies on a leader-based coordination phase to synchronize weights, whereas our proposed solution is fully decentralized, with each load balancer independently adapting weights through local feedback. Nonetheless, both methods share the key objective of distributed QoS maintenance without centralized orchestration.
@@
Zheng et al. [6] extended this research by introducing an auction-based computation-offloading framework under incomplete information. Their system couples distributed multi-armed bandit learning with a heterogeneous auction mechanism to adapt to dynamic network and load variations. The TODAuction-MAB algorithm uses the estimated rewards to decide how to allocate servers to each user through an auction-based mechanism, while the HD-MAB algorithm adds sliding-window updates to handle non-stationary conditions. Despite strong performance and theoretical guarantees, their architecture still relies on a leader node to coordinate distributed auctions and assumes that each server handles at most one user at a time, with the total number of servers exceeding the number of users. These assumptions simplify the offloading model but limit generalization to realistic CC scenarios.
@@
Most of the reviewed approaches primarily focus on minimizing average latency or total delay ([5, 6, 9, 10, 11, 12, 13]), but they rarely formulate QoS as a strict per-client constraint defined by a latency threshold ([4, 7, 14]). Consequently, they aim for global performance improvement rather than ensuring that each client consistently meets its QoS target. Furthermore, these works rarely consider a realistic CC deployment scenario in which service instances are already predeployed, e.g., by an orchestration framework, and can simultaneously serve multiple clients. Among recent studies, there is a clear methodological trend toward modeling routing and offloading as MAB ([5, 6, 15]) or RL ([7]) problems, enabling adaptive decision making under uncertainty. However, many of these methods still rely on partial centralization, such as leader-based coordination in auction mechanisms or explicit inter-node communication for load estimation, and thus fail to achieve full decentralization.
This state-of-the-art review has motivated us to formulate our routing problem as a multi-player multi-armed bandit, enabling adaptive learning of instance selection under uncertainty while ensuring decentralized operation (see Section IV). Instead of relying on one-to-one offloading decisions or coordinated auctions, our approach introduces a fully distributed load balancer that continuously adjusts routing weights based on locally observed QoS feedback. By combining the adaptive exploration–exploitation principles of the MAB framework with dynamic weight adaptation, the proposed QEdgeProxy achieves decentralized and scalable routing that maintains per-client QoS satisfaction and balances load across multiple active instances in real time.
@@ III System Model
  • $K$ load balancers (LBs), indexed by $k\in\mathcal{K}=\{1,\dots,K\}$.
@@
Each LB $k$ serves a client request stream at a known rate $\lambda_{k}\geq 0$, which may vary across LBs. Each client is associated to a single LB responsible for routing its traffic. For each LB $k$, we denote by
@@
  • $W$ is the time window (in time units) over which the QoS requirement is evaluated.
@@
  • $l_{m}^{p}(t)$: request processing time of instance $m$, including queuing and execution delay.
@@
which determines whether the QoS constraint $\ell_{c,m}(t)\leq\tau$ is satisfied or not.
@@ IV MP-MAB Problem Formulation
The runtime load balancing problem is modeled as a multi-player multi-armed bandit (MP-MAB) process with heterogeneous reward, where multiple load balancers (players) concurrently select service instances (arms) to route their incoming requests.
  • Each load balancer (LB) $k\in\mathcal{K}$ maintains a probability distribution (weight vector) $w_{k}(t)$ which determines the fraction of its incoming request rate $\lambda_{k}(t)$ that is routed to each active instance $m\in\mathcal{M}(t)$. We assume that $K\geq M(t)$ holds at any time $t$, reflecting the common CC configuration where service instances are deployed only on a subset of available nodes. By design, our system deploys one LB per node.

### v2

@@ Abstract
As computation shifts from the cloud to the edge to reduce processing latency and network traffic, the resulting Computing Continuum (CC) creates a dynamic environment where meeting strict Quality of Service (QoS) requirements and avoiding service instance overload becomes challenging. Existing methods often prioritize global metrics and overlook per-client QoS, which is crucial for latency-sensitive and reliability-critical applications. We propose QEdgeProxy, a decentralized QoS-aware load balancer that acts as a proxy between IoT devices and service instances in the CC. We formulate the load balancing problem as a Multi-Player Multi-Armed Bandit (MP-MAB) with heterogeneous rewards: Each load balancer autonomously selects service instances to maximize the probability of meeting its clients’ QoS requirements by using Kernel Density Estimation (KDE) to estimate QoS success probabilities. Our load-balancing algorithm also incorporates an adaptive exploration mechanism to recover rapidly from performance shifts and non-stationary conditions. We present a Kubernetes-native QEdgeProxy implementation and evaluate it on an emulated CC testbed deployed on a K3s cluster with realistic network conditions and a latency-sensitive edge-AI workload. Results show that QEdgeProxy significantly outperforms proximity-based and reinforcement-learning baselines in per-client QoS satisfaction, while adapting effectively to load surges and changes in instance availability.
@@ I Introduction
The growing scale of Internet of Things (IoT) deployments has made routing all data to remote cloud servers for processing increasingly inefficient. This centralized model leads to increased latency, congestion, and limited responsiveness for real-time applications. The Computing Continuum (CC) addresses these challenges by distributing computation across IoT, edge, fog, and cloud nodes, enabling data to be processed closer to its source to improve both performance and resource utilization. Despite these advantages, the CC remains a challenging environment as it is highly dynamic: nodes differ in resources, load, availability, and network conditions, while IoT workloads fluctuate over time and space. Ensuring that services deployed across this heterogeneous continuum continue to meet strict Quality of Service (QoS) requirements therefore demands mechanisms that can adapt to changing conditions.
@@
Reliable service delivery in the CC depends on the explicit definition and enforcement of QoS requirements. These requirements specify measurable performance guarantees, typically expressed through Service Level Objectives (SLOs) and Service Level Agreements (SLAs) [1]. SLOs define concrete targets, such as ensuring that a given percentage of requests remain below a latency threshold, while SLAs formalize these guarantees contractually between providers and consumers. QoS requirements in the CC may be expressed along multiple dimensions, including latency, throughput, reliability, availability, privacy, and security. Their importance varies across application domains; for example, industrial control and autonomous systems prioritize ultra-low latency, whereas healthcare applications may emphasize both privacy and availability. Formalizing such QoS constraints provides a foundation for coordinating placement, routing, and adaptation strategies so that services remain within their expected performance boundaries under dynamic and heterogeneous CC conditions.
This work addresses two key challenges that arise once service instances are deployed in the CC across different layers of the continuum: (i) guaranteeing QoS for continuous IoT data processing for each client through adaptive, decentralized request routing under uncertainty, and (ii) maintaining continuous data delivery from multiple independent clients without overloading service instances on resource-constrained devices or compute nodes.
@@
To support QoS-aware routing in dynamic CC conditions, our previous work [2] introduced QEdgeProxy, a lightweight decentralized load balancer (LB) deployed on every CC node. QEdgeProxy acts as an intermediary between clients and service instances, selecting a suitable instance for each client request based on locally observed QoS parameters while continuously adapting its routing decisions as conditions in the CC change. Rather than selecting a single optimal instance that maximizes QoS for each client request, it maintains a pool of instances (QoS pool) capable of meeting the QoS requirements with high probability. This approach enables effective load balancing while preserving per-client QoS guarantees. QEdgeProxy is designed to handle both client-to-service and service-to-service interactions, allowing IoT devices and intermediate services to forward requests through a QEdgeProxy instance deployed on their nearest or hosting node. By maintaining and dynamically updating a QoS pool, QEdgeProxy ensures balanced request distribution and mitigates instance overload while satisfying QoS requirements for each client.
@@
While the high-level architectural design and preliminary proof-of-concept implementation of QEdgeProxy were introduced in [2], the load-balancing logic was limited to constructing QoS pools based on average latency measurements and distributing requests within the pool using simple round-robin scheduling. This paper significantly extends and formalizes the load-balancing mechanism by developing the underlying algorithms for adaptive QoS-pool maintenance and request routing. We formulate the problem as a Multi-Player Multi-Armed Bandit (MP-MAB) with heterogeneous rewards [3] and provide a decentralized learning procedure—built on Kernel Density Estimation (KDE)-based QoS estimation, adaptive exploration, and instance-specific weight updates—that enables QEdgeProxy to operate robustly under dynamic and non-stationary CC conditions. This work provides the theoretical foundation, detailed algorithmic implementation, and a comprehensive evaluation that were not addressed in the initial QEdgeProxy prototype, resulting in a complete and rigorous load-balancing solution.
Unlike most prior work that prioritizes global performance metrics, such as minimizing average latency or total delay [4, 5], or relies on partial coordination among decision makers [6, 7], this paper focuses on guaranteeing per-client QoS requirements through fully decentralized load balancing. By formulating routing as an MP-MAB with heterogeneous and non-stationary rewards, our approach enables each LB to adapt routing decisions based solely on local QoS feedback, without centralized control or inter-node coordination.
@@
  • MP-MAB problem formulation for QoS-aware load balancing. We formulate the load-balancing problem in CC as an MP-MAB with heterogeneous and non-stationary rewards. Unlike classical MAB or offloading models based on single-action selection, our formulation captures weighted request routing, shared service instances across multiple LBs, and strict per-client QoS requirements.
@@
  • Decentralized algorithms for QoS-aware load balancing in CC. We develop a decentralized set of algorithms that solve the proposed MP-MAB formulation and enable QoS-aware load balancing in dynamic CC environments. Each instance of LB autonomously maintains a QoS pool—a set of service instances predicted to meet the required QoS—using a combination of $\varepsilon$-decay exploration and KDE to estimate instance-level QoS success probabilities. Furthermore, it incorporates an adaptive exploration mechanism that reduces exploration as confidence grows, but increases it when QoS degradation or performance shifts are detected, enabling rapid recovery in non-stationary conditions. Requests are routed using a Smooth Weighted Round Robin (SWRR) mechanism parameterized by these instance-level probabilities, ensuring balanced utilization of service instances while preserving per-client QoS guarantees.
@@
  • Kubernetes-native implementation. We implement the algorithms within a Kubernetes-native component (QEdgeProxy), enabling each node to run a decentralized QoS-aware load balancer that maintains QoS pools, performs adaptive weighted routing, and obtains node and instance updates through an MQTT-based informer to reduce the monitoring overhead. This design demonstrates that decentralized QoS-aware load balancing can be seamlessly integrated into existing CC orchestration solutions with minimal configuration.
@@
  • Real-world validation in a distributed CC testbed. We evaluate QEdgeProxy instances deployed across a real K3s cluster emulating a 30-node CC topology with realistic network latencies. Using a latency-sensitive IoT edge-AI workload, we show that our solution outperforms proximity-based and reinforcement-learning baselines, achieving higher per-client QoS satisfaction, improved fairness, robust adaptability to dynamic changes, and low resource overhead suitable for constrained edge devices.
@@
The remainder of this paper is organized as follows. Section II reviews related work on QoS-aware routing, decentralized load balancing, and learning-based decision-making in the CC. Section III introduces the system model, describing both design-time and runtime behavior of the system, and QoS requirements definition. Section IV formalizes the decentralized QoS-aware load balancing problem as an MP-MAB problem with heterogeneous rewards. Section V proposes the decentralized load-balancing algorithms implemented in QEdgeProxy, including QoS-pool construction, KDE-based estimation, $\varepsilon$-decay exploration, and SWRR-based request routing. Section VI details the Kubernetes-native implementation of the QEdgeProxy, and Section VII reports the experimental evaluation conducted on an emulated CC testbed. Finally, Section VIII concludes the paper. In Table I we listed all the symbols used in the paper.
@@
$\mathcal{K}$ Set of load balancers (LB); one per CC node
@@
$\mathcal{M}(t)$ Active service instances
@@
$r_{k,m}(t)$ Reward for LB $k$ routing a request to $m$
@@
$\mu_{k,m}(t)$ Expected QoS success probability for LB $k$ on instance $m$
@@
$\varepsilon(t)$ Exploration factor
@@
* Time-dependent quantities are indexed by discrete time $t$.
@@ II Related work
Fahs and Pierre introduced Proxy-mity [4], a proximity-aware traffic routing system designed as a plugin for Kubernetes. Their goal was to make Kubernetes’ routing aware of geographical distribution by enabling a trade-off between load balancing and latency reduction through a controllable parameter $\alpha$. When $\alpha=0$, traffic is evenly distributed across all service replicas, whereas $\alpha=1$ prioritizes routing to the closest instance. The approach seamlessly integrates with Kubernetes and reduces the request latency by up to 90% in experimental fog setups. However, Proxy-mity considers only network latency as a proximity metric and does not incorporate feedback from actual request processing times. Moreover, it does not allow the explicit definition of QoS constraints such as maximum latency thresholds or success ratios.
Liu and Fang [6] proposed AdeQoS, a decentralized multi-player multi-armed bandit (MP-MAB) algorithm that guarantees per-player QoS satisfaction. Their model assumes that each player (e.g., a load balancer) interacts with a set of arms (service instances) without coordination or observation of others’ actions. Unlike conventional reward maximization, AdeQoS tries to ensure that each player maintains an expected reward above a given QoS threshold, achieving $O(1)$ QoS regret. This approach is closely aligned with our formulation (see Section IV), as both aim to meet per-client QoS constraints rather than maximizing aggregate utility. However, AdeQoS assumes that the number of arms exceeds the number of players and that collisions occur when two players select the same arm, which differs from our model (Section III) where each service instance can handle multiple requests simultaneously. Moreover, AdeQoS relies on a leader-based coordination phase to synchronize weights, whereas our solution is fully decentralized, with each load balancer independently adapting weights through local QoS measurements. Nonetheless, both methods share the key objective of distributed QoS maintenance without centralized orchestration.
@@
Zheng et al. [8] extended this research by introducing an auction-based computation-offloading framework under incomplete information. Their system couples distributed MAB learning with a heterogeneous auction mechanism to adapt to dynamic network and load variations. The TODAuction-MAB algorithm uses the estimated rewards to decide how to allocate servers to each user through an auction-based mechanism, while the HD-MAB algorithm adds sliding-window updates to handle non-stationary conditions. Despite strong performance and theoretical guarantees, their architecture still relies on a leader node to coordinate distributed auctions and assumes that each server handles at most one user at a time, with the total number of servers exceeding the number of users. These assumptions simplify the offloading model but limit generalization to realistic CC scenarios.
@@
Most of the reviewed approaches primarily focus on minimizing average latency or total delay [5, 8, 10, 11, 12, 13, 14], but they rarely formulate QoS as a strict per-client constraint defined by a latency threshold [6, 7, 15]. Consequently, they aim for global performance improvement rather than ensuring that each client consistently meets its QoS target. Furthermore, these works rarely consider a realistic CC deployment scenario in which service instances are already pre-deployed, e.g., by an orchestration framework, and can simultaneously serve multiple clients. Among recent studies, there is a clear methodological trend toward modeling routing and offloading as MAB [5, 8, 16] or RL [7] problems, enabling adaptive decision making under uncertainty. However, many of these methods still rely on partial centralization, such as leader-based coordination in auction mechanisms or explicit inter-node communication for load estimation, and thus fail to achieve full decentralization.
This state-of-the-art review has motivated us to formulate our routing problem as a multi-player multi-armed bandit, enabling adaptive learning of instance selection under uncertainty while ensuring decentralized operation (see Section IV). Instead of relying on one-to-one offloading decisions or coordinated auctions, our approach introduces a decentralized load balancer that continuously adjusts routing weights based on locally observed QoS feedback. By combining the adaptive exploration–exploitation principles of the MAB framework with dynamic weight adaptation, the proposed QEdgeProxy achieves decentralized and scalable routing that maintains per-client QoS satisfaction and balances load across multiple active instances in real time.
@@ III System Model
  • $K$ load balancers, indexed by $k\in\mathcal{K}=\{1,\dots,K\}$.
@@
Each LB $k$ serves a client request stream at a known rate $\lambda_{k}\geq 0$, which may vary across LBs. Each client is associated with a single LB responsible for routing its traffic. For each LB $k$, we denote by
@@
  • $W$ is the time window (in time units) over which the QoS requirements are evaluated.
@@
  • $l_{m}^{p}(t)$: request processing latency of instance $m$, including queuing and execution delay.
@@
which determines whether the QoS constraint $\ell_{k,m}(t)\leq\tau$ is satisfied or not.
@@ IV MP-MAB Problem Formulation
The runtime load-balancing problem is modeled as a multi-player multi-armed bandit (MP-MAB) process with heterogeneous rewards, where multiple load balancers (players) concurrently select service instances (arms) to route their incoming requests.
  • Each load balancer (LB) $k\in\mathcal{K}$ maintains a probability distribution (weight vector) $w_{k}(t)$ which determines the fraction of its incoming request rate $\lambda_{k}(t)$ that is routed to each active instance $m\in\mathcal{M}(t)$. We assume that $K\geq M(t)$ holds at any time $t$, reflecting the common CC configuration where service instances are deployed only on a subset of available nodes and can serve multiple requests simultaneously. By design, our system deploys one LB per node.